How do you solve for the equation (5x+5)^2=100 ? thanks
Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

How do you solve for the equation (5x+5)^2=100 ? thanks
---
You rearrange:
(5x+5)^2 - 100 = 0
---
You rewrite:
(5x+5)^2 - 10^2 = 0
---
You factor:
(5x+5 - 10)(5x+5+10) = 0
---
You simplify:
(5x-5)(5x+15) = 0
---
You solve:
5x = 5 or 5x = -15
x = 1 or x = -3
==========================
Cheers,
Stan H.
==========================

Answer by Alan3354(69443)

(5x+5)^2=100
Take the square root of both sides
5x+5 = ± 10
x+1 = ± 2
x = 1, x = -3